Solve the system of eqations.
x + y = -4
x - y = 10

Answer by checkley77(12844)

x + y = -4
x - y = 10 ADD
---------------
2X=6
X=6/2
X=3 ANS.
3+Y=-4
Y=-4-3
Y=-7 ANS.
PROOF:
3-(-7)=10
3+7=10
10=10